OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

Equation to solve this would be: u ▪ v = ||u|| ||v||cos theta

geerky42 (geerky42):

Yeah. What are you stuck at?

OpenStudy (anonymous):

I'm attempting to solve it through the given equation and through the given examples (I can make a print screen of them as well), but I don't seem to be getting any answers that are provided.

geerky42 (geerky42):

Hmm. Did you note that \(u = \dfrac{2}{3}v\)?

geerky42 (geerky42):

So that mean these vectors are in same direction, right?

OpenStudy (anonymous):

Yes, I noticed that they are. But where did you get 2/3v from?

geerky42 (geerky42):

Well, I used this formula and I got \(\dfrac{u\cdot v}{||u||~||v||} = 1 =\cos\theta\), so I knew these vectors are in same direction.

OpenStudy (anonymous):

Noted.

geerky42 (geerky42):

Hmm actually I am wrong... Sorry, though it is very close to 1...

geerky42 (geerky42):

let's just do calculations

OpenStudy (anonymous):

What is ||u|| and ||v|| in the equation, if you don't mind me asking?

geerky42 (geerky42):

\(||u||\) is magnitude of \(u\) Mathematically if \(u = \langle u_x,u_y\rangle\), then \[\large ||u|| = \sqrt{u_x^2+u_y^2} \]

OpenStudy (anonymous):

That makes a lot more sense now.

geerky42 (geerky42):

Does that clear things up for you?

geerky42 (geerky42):

Great, now you know, so try and recalculate.

OpenStudy (anonymous):

Now when you say \[u \frac{ 2 }{ x }\], what would the u equal?

OpenStudy (anonymous):

Now just realizing that's not a fraction.

geerky42 (geerky42):

About \(u = \dfrac{2}{3}v\)... It's wrong. Because \(\langle2,-4\rangle\neq\dfrac{2}{3}\langle3,-8\rangle = \langle2,-\dfrac{8}{3}\rangle\) I just made mistake that's all...

OpenStudy (anonymous):

Right, but I'm making reference to ||u||.

geerky42 (geerky42):

Exactly what do you mean?

OpenStudy (anonymous):

When you gave the equation \[||u|| =\sqrt{u^2_x...}\] I'm unsure of what the values in the parenthesis would mean.

OpenStudy (anonymous):

*Not parenthesis, square root.I'm sorry.

geerky42 (geerky42):

These are just x and y components of vector